OpenStudy (anonymous):

what is the inverse of f(x)=x/2+3

OpenStudy (abb0t):

Solve for "x" in terms of "y".

OpenStudy (reemii):

write y=x/2+3 and do as abb0t said. when you obtain "x=....stuff with y...", you actually found x='f-inverse'(y)

OpenStudy (reemii):

y=x/2+3 <-> y-3=x/2 <-> etc

OpenStudy (anonymous):

the inverse is y-3=x/2 ???

OpenStudy (whpalmer4):

Your original function is \(f(x) =y = x/2 + 3\)Solve that for x in terms of y: \[y = x/2+3\]\[y-3=x/2\]\[2(y-3)=x\]Now swap x and y: \[2(x-3)=y\]\[f^{-1}(x)=2(x-3)\]

OpenStudy (whpalmer4):

One interesting thing about inverse functions is that they look just like the original function reflected over the line y = x (a line going up and to the right at a 45 degree angle). Here's an example (different functions, but no matter):

OpenStudy (whpalmer4):

The purple line was the \(f(x)\), the olive line was \(f^{-1}(x)\), and the blue is \(y=x\). One way to check your work when doing inverse functions is to make a little graph and make sure the inverse and original functions make that reflection.